OpenStudy (anonymous):

find the y- coordinate of the point of intersection of the lines -2x-3y=1 5x + 2y =1

hartnn (hartnn):

since we need to find y co-ordinate, lets try to eliminate 'x' do you know the method of elimination?

OpenStudy (anonymous):

we eliminate the 2x and the 5x

hartnn (hartnn):

lets start by multiplying 1st equation by 5 multiply each term of -2x-3y=1 by 5 what do u get?

OpenStudy (anonymous):

-10x-15x=5

hartnn (hartnn):

good! now multiply each term of 5x+2y=1 by 2 so that you get 10x for this too!

hartnn (hartnn):

then all you have to do is just add the 2 new equations :)

OpenStudy (anonymous):

10x -5y=10

hartnn (hartnn):

did you multiply each term by 2?

hartnn (hartnn):

5x+2y =1 2*5x + 2*2y = 2*1 ....

OpenStudy (anonymous):

ooops! -7/11

hartnn (hartnn):

correct :)